# Stylebreeder : Exploring and Democratizing Artistic Styles through Text-to-Image Models

### v1

@@ 4.4 Style-based Recommendation
The sheer volume of styles generated by users presents a significant challenge in navigating the landscape of artistic options available. To address this, we showcase a recommendation system that suggests top styles to users based on their previously generated images. This personalized approach is crucial as it helps users discover styles that align with their tastes and past preferences, simplifying their search among a vast array of choices. We formulate our task as a matrix-factorization based recommendation which involves a set of items, where users rate items they have interacted with, thus creating a matrix of user-item ratings. In the context of our problem, users are the creators that generate images, and items are the clusters in which generated images are assigned. We calculate for each user a vector where represents the proportion of images that user has generated within a specific cluster such that . These vectors create a matrix where entries denotes the rating for user for cluster . We employ the SVD algorithm [49, 10] to obtain a prediction for all where is the global average rating. and are the user and item bias terms, respectively. and are the latent factor vectors for item and user , respectively. We minimize the regularized squared error loss and update parameters using Stochastic Gradient Descent [13]. We assess the MAE and RMSE of the predicted ratings against the ground-truth ratings obtaining a mean RMSE of 0.1425 and standard deviation 0.0017 along with a mean MAE of 0.082 and standard deviation 0.001 across all folds. Fig. 5 (d) depicts an example of recommendations for a user based on previously generated styles.
@@ 4.5 Style Atlas for Democratizing Artistic Styles
Since LoRA has become a popular tool for lightweight concept tuning within the community [39], we are including 100 style LoRAs in our style atlas. Fig. 5 (c) displays a screenshot of our platform where users can browse and download LoRA models for styles they find appealing. Additionally, we provide a notebook that enables users to load these downloaded LoRA models and generate personalized images. For further details, please refer to the Supplementary Material.
@@ 5 Limitations and Societal Impact
While our work significantly advances the integration of AI in creative processes, it also presents certain limitations and societal impacts that warrant careful consideration. One of the primary limitations lies in the potential for over-reliance on technology in artistic creation, which could diminish the value and perception of human-driven artistry and creativity. Additionally, the use of AI in art generation raises concerns about copyright and originality, especially when styles closely mimic those of existing artists without clear attribution. From a societal perspective, while our tools aim to democratize art creation, there exists a risk of reinforcing existing biases present in the training data, which could skew the diversity and representation in generated artworks. Moreover, as these technologies become more accessible, there is a potential for misuse, such as creating deceptive images or deepfakes, which could have broader implications for trust and authenticity in digital media. Acknowledging and addressing these challenges is crucial as we continue to explore the intersection of AI and art.
@@ 6 Conclusion
This paper has demonstrated the significant potential of text-to-image diffusion models to explore and catalog the rich tapestry of user-generated artistic styles on the Artbreeder platform. We have successfully identified unique, previously uncharted artistic expressions and demonstrated their application in generating personalized images. Additionally, we demonstrated that a personalized recommendation system enhances user engagement by aligning suggested styles with individual preferences. We also release of an atlas of styles democratizes the access to these innovations, allowing users to experiment with and adopt new artistic expressions. This work not only advances the technological capabilities of AI in the field of art but also contributes to a more inclusive and diverse artistic community. Our dataset offers numerous avenues for further exploration, such as refining the effectiveness of text prompts through iterative adjustments, studying trends in art over time, recommending styles based on both image and textual content, developing a search system for the generated images, and exploring explainable creativity [29].
@@ Appendix C Croissant Metadata
Appendix D Style Atlas
@@ Appendix E More Comparisons with Other Datasets
Tab 3 features several prominent image synthesis benchmarks. Certain benchmarks focus on single-category image datasets for unconditional image synthesis evaluation, such face datasets like CelebA [28] and FFHQ [23] (human faces) and MetFaces [22] (face artworks). There are also multi-class datasets like STL-10 [6] and ImageNet [7], but these datasets predominantly consist of photographic images with a limited representation of artwork. ArtBench-10 [27] is primarily composed of artwork, however, with only 10 classes, it heavily skews toward artwork of North American, European, and East Asian origin and fails to encompass digital and modern art. DiffusionDB [55] and TWIGMA [5] present AI-generated art datasets, but suffer from a lack of model diversity and original generation prompts, respectively.
@@
| Ours | (512, 512) | (1280, 986) | 6,818,217 | AI artworks |
Appendix F Artist names
On platforms such as Artbreeder, it is common to use artist names in text-prompts to indicate a particular style. We used a popular NER library [52] to extract artist names from the text prompts. Top 20 artist names are provided in Table 4. We provided an additional column in our dataset to indicate potential artist names in the text-prompts. Moreover, we provide a Google form on our website if any artist whose name is used would like to opt-out from our dataset.
@@ Appendix F Artist names
| Adriana Lima | 84967 |
Appendix G Datasheet
@@
As artists worldwide are increasingly leveraging text-to-image diffusion models to create artworks spanning diverse set of styles, there is a pressing need uncover and promote unique artistic expressions. This will enable users to discover unique artistic styles, generate personalized content using those styles, and get recommendations about styles they would be interested. STYLEBREEDER project was created to address these questions, and to the best of our knowledge, has the largest userbase (95K) spanning to 6.8M images with 1.8M unique text prompts.
@@
Each instance consists of an image generated by a text-to-image diffusion model (such as SD-XL or ControlNet) and the original text prompt used for generating the image, as well as original metadata and expanded features created by us. The metadata originally collected from the Artbreeder website are: Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale. We also offer further metadata like Cluster ID, along with scores for Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack computed with [15, 16]. We also provide potential artist names that are used in the text prompt such as “Ilya Kuvshinov”.
@@
6,818,217 images created by 95,479 unique users, with 1.8M unique text-prompts.
@@
No tests were run in order to determine representativeness of the dataset. However, our dataset consists of all publicly generated images on Artbreeder Platform’s Collage tool, from July 2022-May 2024. Out of this dataset, we provide a 2M version which includes all data as well as the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder. We provide all features except images for the full 6.8M dataset on the same website, but provide a script to scrape the 6.8M images from Artbreeder platform. If requested by the reviewer committee at NeurIPS, we will try our best to find a platform that can accommodate the full image data.
@@
Each instance contains a unique image ID, corresponding meta and expanded features including Positive Prompt, Negative Prompt, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale, Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names. We also provide the original image files along with 2M subset https://huggingface.co/datasets/stylebreeder/stylebreeder, and a script to download 6.8M version.
@@
For each image, we provide an extended set of labels including text prompt, model parameters, and other features as discussed above.
@@
There is no missing data.
@@
Each image is connected to an anonymized user ID. By filtering the dataset by user ID, one can infer relationship between different samples, where those samples are created by the same user.
@@
rationale behind them.
@@
No, all images and related metadata are collected directly from Artbreeder platform.
@@
We provide a 2M version which includes all data as well as the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder. We provide all features except images for the full 6.8M dataset on the same website, but provide a script to scrape the 6.8M images from Artbreeder platform. If requested by the reviewer committee at NeurIPS, we will try our best to find a platform that can accommodate the full image data.
@@
The dataset is collected from user’s publicly available content, however since there is no control over what type of text prompts users can input, it is possible that some confidential information might be included in the text prompts, however we expect such cases to be rare.
@@
Similar to other datasets with AI-generated images like DiffusionDB, TWIGMA, ours contain a substantial amount of potentially NSFW content both in terms of images and text prompts. This observation correlates with recent studies that highlight a significant increase in NSFW content generation by online communities [36, 41]. We provide an analysis on NSFW and other features such as Toxicity, Severe Toxicity, Identity Attack, Obscene, Insult and Threat in Fig 3 and Section 3.4 in the main paper. Moreover, to help researchers to filter out potentially unsafe content we provide these scores in our dataset.
@@
We share anonymized user IDs and text prompts may have celebrity or artist names in the text prompts.
@@
Text prompts may have celebrity or artist names in the text prompts. We provide a Google form for reporting harmful or inappropriate images and prompts, as well as allowing artists to opt-out in case their names are used in the text prompts.
@@
We do not collect any user-related data except anonymized user ID.
@@
The images, prompt and other metadata including including Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters including Model Type, Seed, Step, and CFG Scale are collected from Artbreeder website. Moreover we derived additional features such as Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names [15, 16, 53].
@@
All data was collected from April-May 2024, The images generated between July 2022-May 2024. The dataset first published with this NeurIPS submission, June 10, 2024.
@@
The image IDs and original metadata is directly pulled from the database by one of the authors Joel Simon, owner of Artbreeder platform. A Python script was used to scrape the images, and infer additional features such as Cluster ID, or NSFW scores.
@@
The metadata and image IDs were provided directly by Artbreeder via their database. Collecting the corresponding images took a total of 78 hours, estimated bandwidth Costs: $315, storage Costs: $80.5 and Energy Costs: $3.9 (assuming an electricity cost of $0.10 per kWh to run a server for 78 hours).
@@
We collected all data from Artbreeder’s Collage tool from July 2022-May 2024, so this is not a subset.
@@
Students collected the data as a part of their graduate research assistantship by Virginia Tech or ETH Zurich.
@@
No, by Artbreeder’s contract every user accepts the agreement that their data is licensed as CC0 (Public Domain). Exact language can be seen from: https://www.artbreeder.com/terms.pdf.
@@
Artbreeder’s contract every user accepts the agreement that their data is licensed as CC0 (Public Domain), hence we did not ask for additional consent.
@@
Users will have the option to report any harmful, unsafe and sensitive content via our Google form on our website: http://stylebreeder.github.io. Also, artists whose names are used in the text prompts has the option to opt-out via the same form.
@@
We used a library to infer various scores such as NSFW, Toxicity, Insult, Threat, Identity Attack scores using Detoxify library [15]. Image NSFW scores are computed using NSFW-Detector library [16]. To infer artist names, we used Named Entity Recognition via [53].
@@
Since we collected full data as opposed to a subset, we do not expect any unfair treatment of individuals or groups. Moreover, since the dataset is already in public domain, we do not expect any legal or financial risks. However, as discussed in our paper, some of the images and text prompts reflect NSFW content, and researchers should use the associated scores provided by us to filter out unsafe data based on their tasks.
@@
All tasks utilize this dataset should follow CC0 Public domain licensing rules.
@@
We share the dataset, code and other information on our website; http://stylebreeder.github.io.
@@
All images generated on Artbreeder platform are under the CC0 Public Domain License.
@@
Authors of this paper are responsible for supporting, hosting and maintaining the dataset. We also provide a Google form for users to report any potential issue about the dataset.
@@
Authors of this paper will monitor the Google form where users can report any potential issues. We will update our datasets biweekly, and announce the updates on our website: http://stylebreeder.github.io.
@@
Other researchers are welcome to extend or contribute to our dataset, they can either send a pull request to our github repo (http://github.com/stylebreeder) or contact authors on Google form.

### v2

@@ 4.4 Style-based Recommendation
The sheer volume of styles generated by users presents a significant challenge in navigating the landscape of artistic options available. To address this, we showcase a recommendation system that suggests top styles to users based on their previously generated images. This personalized approach is crucial as it helps users discover styles that align with their tastes and past preferences, simplifying their search among a vast array of choices. We formulate our task as a matrix-factorization-based recommendation, which involves a set of items where users rate items they have interacted with, thus creating a matrix of user-item ratings. In the context of our problem, users are the creators who generate images, and items are the clusters in which generated images are assigned. We calculate for each user a vector where represents the proportion of images that the user has generated within a specific cluster such that . These vectors create a matrix where entries denote the rating for user for cluster . We employ the SVD algorithm [51, 10] to obtain a prediction for all where is the global average rating. and are the user and item bias terms, respectively. and are the latent factor vectors for item and user , respectively. We minimize the regularized squared error loss and update parameters using Stochastic Gradient Descent [13]. We assess the MAE and RMSE of the predicted ratings against the ground-truth ratings, obtaining a mean RMSE of 0.1425 and a standard deviation of 0.0017, along with a mean MAE of 0.082 and a standard deviation of 0.001 across all folds. Fig. 5 (d) depicts an example of recommendations for a user based on previously generated styles.
@@ 4.5 Style Atlas for Democratizing Artistic Styles
Since LoRA has become a popular tool for lightweight concept tuning within the community [41], we provide 100 style LoRAs in our Style Atlas platform. Fig. 5 (c) displays a screenshot of our platform where users can browse and download LoRA models for appealing styles. Additionally, we provide a notebook that enables users to load these downloaded LoRA models and generate personalized images. For further details, please refer to the Supplementary Material.
@@ 5 Limitations and Societal Impact
While our work significantly advances the integration of AI in creative processes, it also presents certain limitations and societal impacts that warrant careful consideration. One of the limitations lies in the potential for over-reliance on technology in artistic creation, which could diminish the value and perception of human-driven artistry and creativity. Additionally, the use of AI in art generation raises concerns about copyright and originality, especially when styles closely mimic those of existing artists without clear attribution. From a societal perspective, while our tools aim to democratize art creation, there exists a risk of reinforcing existing biases present in the training data, which could skew the diversity and representation of generated artworks. Moreover, as these technologies become more accessible, there is a potential for misuse, such as creating deceptive images or deepfakes, which could have broader implications for trust and authenticity in digital media. Acknowledging these challenges is crucial as we continue to explore the intersection of AI and art.
@@ 6 Conclusion
This paper has demonstrated the significant potential of text-to-image diffusion models to explore and catalog the rich tapestry of user-generated artistic styles on the Artbreeder platform. We successfully identify unique, previously uncharted artistic expressions and demonstrate their application in generating personalized images. Additionally, we demonstrate that a personalized recommendation system enhances user engagement by aligning suggested styles with individual preferences. We also release the Style Atlas platform that democratizes access to these innovations, allowing users to experiment with and adopt new artistic expressions. This work not only advances the technological capabilities of AI in the field of art but also contributes to a more inclusive and diverse artistic community. Our dataset offers numerous avenues for further exploration, such as refining the effectiveness of text prompts through iterative adjustments, studying trends in art over time, recommending styles based on both image and textual content, developing a search system for the generated images, and exploring explainable creativity [30].
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] See Section 5.
Did you discuss any potential negative societal impacts of your work? [Yes] See Section 5.
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Please see Section 4.1.
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We provide standard deviations in Tab. 2 where both CLIP and DINO scores were computed on 20K images each.
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Please see Section 4.1.
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] We cited the Artbreeder platform for curating the assets.
Did you mention the license of the assets? [Yes] Our dataset and code are released with a CC0 Public Domain license.
Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide our code and dataset as a URL https://stylebreeder.github.io.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Users on the Artbreeder platform agree to release their data under CC0 Public Domain license, as discussed in Section 3.
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] All user IDs in our dataset have been anonymized to ensure users’ privacy. We also extensively analyze offensive content, see Section 3.4.
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ Appendix C Croissant Metadata
Appendix D Style Atlas Platform
@@ Appendix E More Comparisons with Other Datasets
Table 3 features several prominent image synthesis benchmarks. Certain benchmarks focus on single-category image datasets for unconditional image synthesis evaluation, such as face datasets like CelebA [29] and FFHQ [24] (human faces) and MetFaces [23] (face artworks). There are also multi-class datasets like STL-10 [6] and ImageNet [7], but these datasets predominantly consist of photographic images with a limited artwork representation. ArtBench-10 [28] is primarily composed of artwork. However, with only ten classes, it heavily skews toward artwork of North American, European, and East Asian origin and fails to encompass digital and modern art. TWIGMA [5] presents an AI-generated art dataset but suffers from a lack of original generation prompts and a significantly smaller volume of images. The main distinction between our dataset and other AI-generated datasets of similar scale, such as DiffusionDB [57], lies in the duration over which images were generated. While DiffusionDB covers images generated during a 12-day period in August 2022, STYLEBREEDER extends across a much longer time frame, spanning 18 months from July 2022 to May 2024. This extensive duration provides a significant advantage for in-depth studies into the evolution and dynamics of visual trends, artistic styles, and thematic content. By covering a broader range of temporal variations, our dataset allows for a more detailed analysis of how generative models respond to changing cultural or seasonal influences over time.
@@
| STYLEBREEDER (Ours) | (512, 512) | (1280, 986) | 6,818,217 | AI artworks |
Appendix F Temporal Trends in Seasonal Content Generation
Our dataset offers potential for deeper exploration of temporal patterns, particularly how seasonal variations influence content generation throughout the year. For instance, a preliminary analysis of the week preceding Halloween (October 24–31) revealed a marked increase in keywords such as Halloween, scary, costume, and pumpkin compared to other periods. Further examination of these seasonal trends could yield valuable insights, which we leave as an avenue for future work.
Appendix G Detailed Analysis on Clustering
We leverage the silhouette score to determine the optimal number of clusters at (as can be seen from Tab. 4, we also try 20000 clusters, but it did not offer a significant deviation in silhouette score compared to 10000 clusters). This number of clusters provides the most meaningful and distinct categorization of our data, which is then used for KMeans++ clustering.
| # Clusters | Silhouette Score |
| 50 | 0.032 |
| 100 | 0.043 |
| 500 | 0.054 |
| 1000 | 0.064 |
| 2000 | 0.078 |
| 5000 | 0.087 |
| 10000 | 0.110 |
| 20000 | 0.111 |
We aim to capture artistic styles within each of these clusters using CSD [46] and hypothesize that the majority of the clusters focus on individual artists with unique styles or groups of artists who have similar styles. Further analysis of the number of assigned clusters to unique artist names reveals interesting distribution patterns. We examine the dominance of individual artists within clusters, inspecting the number of clusters where the top contributing artists represented a significant portion of the data points, and observe the following:
1551 clusters are dominated by a single artist: This means that in these clusters, over 50% of the data points belong to a single artist, highlighting a strong association between the cluster and that artist’s distinct style.
2345 clusters are dominated by two artists: This suggests that these clusters capture stylistic similarities between two artists, potentially representing shared influences, overlapping techniques, or a broader stylistic movement encompassing both artists.
1467 clusters are dominated by the three artists: This further expands the scope of shared stylistic traits, indicating potential sub-genres or broader artistic trends encompassing a small group of artists.
884 clusters are dominated by four artists: This reinforces the trend of clusters capturing shared stylistic qualities among a small group of artists, suggesting the presence of broader artistic movements or schools of thought.
This reveals a fascinating dynamic between individual artistic styles and broader trends. While a significant portion of clusters (1551 or 15.5% of the total 10000 clusters) strongly represent individual artists, a larger portion (almost 52% when considering clusters dominated by the top two, three, or four artists) suggests the presence of shared stylistic traits and broader artistic movements.
Appendix H Identifying Artistic Influences in Text Prompts
On platforms such as Artbreeder, it is common to use artist names in text prompts to indicate a particular style. We used a popular NER library [54] to extract artist names from the text prompts. The top 18 artist names are provided in Tab. 5. We provided another column in our dataset to indicate potential artist names in the text prompts. Moreover, we provide a Google form on our website if any artist whose name is used would like to opt-out from our dataset.
@@ Appendix H Identifying Artistic Influences in Text Prompts
Appendix I Copyright Infringement
One of the primary applications of this large-scale dataset of generated content is to understand how much-generated art is a function of an original artist’s work. This has a significant impact on the copyright infringement of T2I models. We apply the method from [34], which performs artistic style classification by determining a particular artist’s unique style through a reference dataset of 372 artists’ works. This enables us to recognize if identified styles reappear within images in our dataset. Using this successfully identified all 372 artists supported, resulting in 688K images within our dataset. The top artists discovered using this dataset are highlighted in Tab. 6. This demonstrates that our dataset can serve as a valuable resource for scaling up the mentioned method, thereby improving the robustness of artist classification across the 372 artists. Additionally, given the extensive coverage of artists in our dataset, this method could be extended to cover a much broader range of artists, offering valuable insights for addressing copyright infringement issues.
| Artist Name | Accuracy | # Samples in STYLEBREEDER |
| Alphonse Mucha | 42% | 209357 |
| Ivan Aivazovsky | 39% | 50177 |
| Francis Bacon | 28% | 5565 |
| Claude Monet | 28% | 8210 |
| Vincent Van Gogh | 28% | 14671 |
| Hieronymus Bosch | 27% | 8098 |
| Frank Stella | 23% | 9254 |
| John William Waterhouse | 22% | 60683 |
| Egon Schiele | 21% | 11003 |
| Dan Witz | 21% | 10964 |
Appendix J Datasheet
@@ Appendix I Copyright Infringement
As artists worldwide are increasingly leveraging text-to-image diffusion models to create artworks spanning diverse sets of styles, there is a pressing need to uncover and promote unique artistic expressions. This will enable users to discover unique artistic styles, generate personalized content using those styles, and get recommendations about styles they would be interested. STYLEBREEDER project was created to address these questions, and to the best of our knowledge, has the largest userbase (95K) spanning to 6.8M images with 1.8M unique text prompts.
@@
Each instance consists of an image generated by a text-to-image diffusion model (such as SD-XL or ControlNet) and the original text prompt used for generating the image, as well as original metadata and expanded features created by us. The metadata originally collected from the Artbreeder website are Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, Image Size (height, width), and model-related hyperparameters, including Model Type, Seed, Step, and CFG Scale. We also offer further metadata like Cluster ID, along with scores for Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, and Identity Attack computed with [15, 16]. We also provide potential artist names that are used in the text prompt, such as “Ilya Kuvshinov”.
@@
6,818,217 images were created by 95,479 unique users, with 1.8M unique text prompts.
@@
No tests were run to determine the representativeness of the dataset. However, our dataset consists of all publicly generated images on the Artbreeder Platform’s Collage tool from July 2022 to May 2024. We provide a 2M version which includes all data and the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder. Due to resource constraints, we provide a script to download the full data on our website: https://huggingface.co/datasets/stylebreeder/stylebreeder. The full data is divided into a 5M version and a 6.8M version. The 5M version includes images filtered by LLavaGuard [17] to remove potentially unsafe content. The 6.8M version includes both safe and unsafe text prompts and will be shared with researchers to promote research in safety and ethics, and will be shared upon receiving information regarding intended usage via this Google Form: https://forms.gle/FtQ44igxpbgz6LHe8.
@@
Each instance contains a unique image ID, corresponding meta, and expanded features, including Positive Prompt, Negative Prompt, anonymized UserID, Timestamp, Image Size (height, width), and model-related hyperparameters, including Model Type, Seed, Step, and CFG Scale, Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names. We also provide the original image files along with 2M subset https://huggingface.co/datasets/stylebreeder/stylebreeder, and a script to download the 5M version.
@@
For each image, we provide an extended set of labels, including text prompt, model parameters, and other features as discussed above.
@@
The public version of the dataset is filtered by LLavaGuard [17] to remove potentially unsafe content.
@@
Each image is connected to an anonymized user ID. By filtering the dataset by user ID, one can infer the relationship between different samples where the same user creates those samples.
@@
the rationale behind them.
@@
No, all images and related metadata are collected directly from the Artbreeder platform.
@@
We provide a 2M version which includes all data and the original images, hosted on a publicly available platform: https://huggingface.co/datasets/stylebreeder/stylebreeder.
@@
The dataset is collected from users’ publicly available content; however, since there is no control over what type of text prompts users can input, it is possible that some confidential information might be included in the text prompts; however, we expect such cases to be rare.
@@
Similar to other datasets with AI-generated images like DiffusionDB and TWIGMA, ours contains a substantial amount of potentially NSFW content regarding images and text prompts. This observation correlates with recent studies highlighting a significant increase in NSFW content generation by online communities [38, 43]. We provide an analysis on NSFW and other features such as Toxicity, Severe Toxicity, Identity Attack, Obscene, Insult, and Threat in Fig. 4 and Sec. 3.4 in the main paper. Moreover, to help researchers to filter out potentially unsafe content we provide these scores in our dataset.
@@
We share anonymized user IDs, and text prompts may have celebrity or artist names in the text prompts.
@@
Text prompts may have celebrity or artist names in them. We provide a Google form for reporting harmful or inappropriate images and prompts, as well as allowing artists to opt-out in case their names are used in the text prompts.
@@
We do not collect any user-related data except anonymized user IDs.
@@
The images, prompt, and other metadata, including Positive Prompt, Negative Prompt, anonymized ImageID, anonymized UserID, Timestamp, and Image Size (height, width), and model-related hyperparameters, including Model Type, Seed, Step, and CFG Scale, are collected from Artbreeder website. Moreover, we derived additional features such as Cluster ID, Prompt NSFW, Image NSFW, Toxicity, Insult, Threat, Identity Attack scores, Artist names [15, 16, 55].
@@
All data was collected from April-May 2024, The images generated between July 2022-May 2024. The dataset was first published with this NeurIPS submission on June 10, 2024.
@@
The image IDs and original metadata are directly pulled from the database by one of the authors, Joel Simon, owner of the Artbreeder platform. A Python script was used to scrape the images and infer additional features such as Cluster ID or NSFW scores.
@@
Artbreeder provided the metadata and image IDs directly via their database. Collecting the corresponding images took a total of 78 hours, estimated bandwidth Costs: $315, storage Costs: $80.5, and Energy Costs: $3.9 (assuming an electricity cost of $0.10 per kWh to run a server for 78 hours).
@@
We collected all data from Artbreeder’s Collage tool from July 2022 to May 2024, so this is not a subset.
@@
Students collected the data as a part of their graduate research assistantship at Virginia Tech or ETH Zurich.
@@
No, according to Artbreeder’s contract, every user accepts the agreement that their data is licensed as CC0 (Public Domain). The exact language can be seen from: https://www.artbreeder.com/terms.pdf.
@@
In Artbreeder’s contract, every user accepts the agreement that their data is licensed as CC0 (Public Domain). Hence, we did not ask for additional consent.
@@
Users will have the option to report any harmful, unsafe and sensitive content via our Google form on our website: http://stylebreeder.github.io. Also, artists whose names are used in the text prompts can opt-out via the same form.
@@
We used a library to infer various scores such as NSFW, Toxicity, Insult, Threat, and Identity Attack scores using Detoxify library [15]. Image NSFW scores are computed using NSFW-Detector library [16]. To infer artist names, we used Named Entity Recognition via [55].
@@
Since we collected full data instead of a subset, we do not expect any unfair treatment of individuals or groups. Moreover, since the dataset is already in the public domain, we do not expect any legal or financial risks. However, as discussed in our paper, some of the images and text prompts reflect NSFW content, and researchers should use the associated scores we provided to filter out unsafe data based on their tasks.
@@
All tasks utilizing this dataset should follow CC0 Public domain licensing rules.
@@
We share the dataset, code, and other information on our website: http://stylebreeder.github.io.
@@
All images generated on the Artbreeder platform are under the CC0 Public Domain License.
@@
The authors of this paper are responsible for supporting, hosting, and maintaining the dataset. We also provide a Google form for users to report any potential issue about the dataset.
@@
Authors of this paper will monitor the Google form where users can report any potential issues. We will announce the updates on our website: http://stylebreeder.github.io.
@@
Other researchers are welcome to extend or contribute to our dataset. They can either send a pull request to our GitHub repository (http://github.com/stylebreeder) or contact authors on Google form.